OpenStudy (anonymous):

solve for x and y [x y] - [-y x]= [4 6] [-y x] [-y x] [-4 6] I have no clue what to do with this problem.

OpenStudy (anonymous):

this is a matrix problem

OpenStudy (anonymous):

U need to solve the left first then make an equation with the right...

OpenStudy (anonymous):

thats the problem when i was junior school...

OpenStudy (anonymous):

can you show me an example?

OpenStudy (anonymous):

( x^2 + y^2 )- (-xy-xy) = 48 x^2 + y^2 +2xy = 48

OpenStudy (anonymous):

that be (x+y)^2 = 48 (x+y) = \[\sqrt{48}\]

OpenStudy (anonymous):

Ok I see

OpenStudy (anonymous):

thank you